Were do the graphs of the lines y=3 and 4x-3y=6 intersect?
Answer by reviewermath(1029)

Substitute y=3 in 4x-3y=6, the result is
4x-3(3)=6
4x-9=6
4x=15
x=15/4 and y=3
Answer: (15/4,3)